How many molecules (not moles ) of NH 3 are produced from 5.29 * 10 ^ - 4 g of H 2 ?
1
Expert's answer
2020-10-05T14:24:20-0400

Question #136850

How many molecules (not moles) of NH 3 are produced from 5.29 * 10 ^ - 4 g of H 2 ?

Answer

N2+3H2 →2NH3

Mol of H2 used= 5.29×10-4÷2

                           = 0.0002645 mol

Mol of NH3 produced = 0.0002645 mol× 2/3

                                      = 0.000176333

1 mol of NH3 → 6.022×1023 molecules

0.000176333 mol →? molecules

 = 0.000176333× 6.022×1023 molecules

 = 1.061879333×1020 molecules

 = 1.06×1020 molecules